# Adaptive Flip Graph Algorithm for Matrix Multiplication

### 28 Dec 2023 v1

@@ Abstract
This study proposes the “adaptive flip graph algorithm”, which combines adaptive searches with the flip graph algorithm for finding fast and efficient methods for matrix multiplication. The adaptive flip graph algorithm addresses the inherent limitations of exploration and inefficient search encountered in the original flip graph algorithm, particularly when dealing with large matrix multiplication. For the limitation of exploration, the proposed algorithm adaptively transitions over the flip graph, introducing a flexibility that does not strictly reduce the number of multiplications. Concerning the issue of inefficient search in large instances, the proposed algorithm adaptively constraints the search range instead of relying on a completely random search, facilitating more effective exploration. Numerical experimental results demonstrate the effectiveness of the adaptive flip graph algorithm, showing a reduction in the number of multiplications for a $4\times 5$ matrix multiplied by a $5\times 5$ matrix from $76$ to $73$, and that from $95$ to $94$ for a $5\times 5$ matrix multiplied by another $5\times 5$ matrix. These results are obtained in characteristic two.
1 Introduction
@@
Efficient matrix multiplication is crucial for accelerating various computational processes, influencing overall computational speed significantly. Therefore, numerous studies have been undertaken to reduce the computational cost from various perspectives. The groundbreaking work of Strassen demonstrated that two $2\times 2$ matrices can be multiplied with only $7$ multiplications (Strassen, 1969), leading to extensive studies on the complexity of matrix multiplication and the discovery of optimal multiplications for matrices beyond $2\times 2$. For the complexity of matrix multiplication, recent developments continue to aim at finding the upper bound on the matrix multiplication exponent $\omega$. The current best-known bound is $\omega<2.371552$ (Williams et al., 2023), improving the previous best bound of $\omega<2.371866$ (Duan et al., 2022). For optimal multiplication methods, notable progress has been made, reducing the multiplication of two $3\times 3$ matrices to $23$ (Laderman, 1976), and the multiplication of two $4\times 4$ matrices have been reduced to $49$ using recursively applying Strassen’s algorithm (Strassen, 1969). Recently, a reinforcement learning approach has further refined the multiplication of two $4\times 4$ matrices to $47$ (Fawzi et al., 2022) and improved the best-known result for two $5\times 5$ matrices from $98$ to $96$ multiplications. Subsequently, focusing on the symmetry in matrix multiplications, a topology called flip graph has been introduced (Kauers and Moosbauer, 2023). By randomly repeating the local transition of vertices over the flip graph, the best-known result for two $5\times 5$ matrices has been improved from $96$ to $95$ multiplications (Kauers and Moosbauer, 2022, 2023).
@@
In this study, we first identify two practical issues confronting the flip graph algorithm. The first issue is that the local transition can encounter troublesome “states” which limit further exploration and prevent updating the best number of multiplications. The second issue is that a random search on the flip graph becomes impractical as the matrix size increases within a reasonable computation time. In fact, without setting the results obtained by the reinforcement learning to the initial condition, it would take too long to search from the initial condition and the best number of multiplications would not be reached (Kauers and Moosbauer, 2023). To address these issues, this study proposes an “adaptive flip graph algorithm”. For the first issue, when encountering troublesome states, the algorithm adaptively transitions vertices in a direction that does not necessarily reduce the number of matrix multiplications, facilitating an escape from the states. For the second issue, instead of conducting a completely random search, the algorithm adaptively constraints the search range, enabling faster exploration. Experiments show that the adaptive flip graph algorithm reduces the number of multiplication for a $4\times 5$ matrix and a $5\times 5$ matrix from $76$ to $73$ and the number of multiplication of two $5\times 5$ matrices from $95$ to $94$, in characteristic 2.
1.1 Problem Formulation of Matrix Multiplication
@@
Definition 1 (Matrix Multiplication Tensor).
@@
$\mathcal{M}_{n,m,p}=\sum_{i,j,k=1}^{n,m,p}a_{i,j}\otimes b_{j,k}\otimes c_{k,i}\in K^{n\times m}\otimes K^{m\times p}\otimes K^{p\times n}.$ (1)
@@
Example 1 (Standard Classical Matrix Multiplication Algorithm for two $2\times 2$ matrices).
$\mathcal{M}_{2,2,2}$ $=a_{1,1}\otimes b_{1,1}\otimes c_{1,1}$
$+a_{1,2}\otimes b_{2,1}\otimes c_{1,1}$
$+a_{1,1}\otimes b_{1,2}\otimes c_{2,1}$
$+a_{1,2}\otimes b_{2,2}\otimes c_{2,1}$
$+a_{2,1}\otimes b_{1,1}\otimes c_{1,2}$
$+a_{2,2}\otimes b_{2,1}\otimes c_{1,2}$
$+a_{2,1}\otimes b_{1,2}\otimes c_{2,2}$
$+a_{2,2}\otimes b_{2,2}\otimes c_{2,2}.$
@@
Example 2 (Strassen’s Algorithm).
@@
In the following explanation, we define the following for these rank-one tensors.
Definition 2 (Matrix Multiplication Scheme).
We call a finite set $S$ of rank-one tensors whose sum is $\mathcal{M}_{n,m,p}$ $(n,m,p)$-matrix multiplication scheme and $|S|$ is the rank of the scheme.
1.2 Flip Graph Algorithm
A graph, known as the flip graph, has been introduced to represent the transformation relationships between matrix multiplication schemes. A search method on this graph has been proposed to reduce the rank of the scheme (Kauers and Moosbauer, 2023). This algorithm, which is the basis of the one employed in our study, is reviewed in this subsection.
Definition 3 (Flip).
@@ 1.2 Flip Graph Algorithm
Definition 4 (Reduction).
[…]
  1. 1.
@@
  2. 2.
[…]
@@
Definition 5 (Flip Graph).
@@
3:         $S\leftarrow$ Flip of $S$ by Definition 3
@@
8:         $S\leftarrow$ Reduction of $S$ by Definition 4
@@
2 Limitation and Practical Issues of Flip Graph Algorithm
@@
2.1 Non-Reduction States in Principle
Numerical experiments have revealed a phenomenon where, for the $(4,4,4)$-matrix multiplication scheme, the rank of the scheme does not decrease during the search, even with increasing the number of iterations. This phenomenon is considered to be due to the absence of vertices that can reduce the rank among the vertices accessible by the flip algorithm. Such a state, called a “non-reduction state", becomes more common as the matrix size increases.
2.2 Practical Non-Reducibility for Large Matrix Multiplication
In random search experiments for the $(5,5,5)$-matrix multiplication scheme, we observed another phenomenon: the rank does not reduce at all from the initial scheme, which is the standard classical matrix multiplication algorithm as shown in Example 1. Notably, this phenomenon is not observed for a $4\times 4$ matrix or lower and only appears for order $5$ or larger matrices. This observation implies that as the size of the tensor increases, reaching reducible vertices by the random search becomes more challenging.
The following argument shows that this issue is distinct from the non-reduction states described above. Suppose we take the standard classical matrix multiplication algorithm as in Example 1 for the initial scheme. If the matrix is decomposed into submatrices, Strassen’s algorithm as in Example 2 can eventually be applied to a $2\times 2$ submatrix, meaning that there is a reducible vertex in the neighborhood of the initial scheme. This consideration shows that this initial scheme is not in non-reduction states. Even then, it is challenging in a random search to start the rank reduction from the initial scheme.
3 Adaptive Flip Graph Algorithm
@@
3.1 Plus Transition
To address the former issue, we propose the plus transition, which introduces a transition that increases the rank of the scheme. This provides a strategy to escape from the non-reduction state described above, allowing the exploration of schemes that are inaccessible by the flip graph algorithm. The procedure for increasing the rank is defined as a "Plus" transformation.
Definition 6 (Plus).
For a scheme $S$ the $\alpha^{(i)}\neq\alpha^{(j)},\beta^{(i)}\neq\beta^{(j)},\gamma^{(i)}\neq\gamma^{(j)}$ are satisfied, the “Plus" transformation from $S$ to $S^{\prime}$ is defined as follows:
@@ 3.1 Plus Transition
$\rightarrow\alpha^{(i)}\otimes\beta^{(i)}\otimes\gamma^{(i)}+\alpha^{(i)}\otimes\beta^{(j)}\otimes\gamma^{(j)}+(\alpha^{(j)}-\alpha^{(i)})\otimes\beta^{(j)}\otimes\gamma^{(j)}$
$\rightarrow\alpha^{(i)}\otimes(\beta^{(i)}+\beta^{(j)})\otimes\gamma^{(i)}+\alpha^{(i)}\otimes\beta^{(j)}\otimes(\gamma^{(j)}-\gamma^{(i)})+(\alpha^{(j)}-\alpha^{(i)})\otimes\beta^{(j)}\otimes\gamma^{(j)}.$
@@
12:         $S_{t}\leftarrow$ Plus $S_{t}$ by Definition 6
@@ 3.2 Edge Constraints
4 Results

### 16 Mar 2024 v2

@@ Abstract.
This study proposes the “adaptive flip graph algorithm”, which combines adaptive searches with the flip graph algorithm for finding fast and efficient methods for matrix multiplication. The adaptive flip graph algorithm addresses the inherent limitations of exploration and inefficient search encountered in the original flip graph algorithm, particularly when dealing with large matrix multiplication. For the limitation of exploration, the proposed algorithm adaptively transitions over the flip graph, introducing a flexibility that does not strictly reduce the number of multiplications. Concerning the issue of inefficient search in large instances, the proposed algorithm adaptively constraints the search range instead of relying on a completely random search, facilitating more effective exploration. In particular, the formal proof is provided that the introduction of plus transitions in the proposed algorithm ensures the connectivity of any node in the flip graph, representing a method of matrix multiplication. Numerical experimental results demonstrate the effectiveness of the adaptive flip graph algorithm, showing a reduction in the number of multiplications for a $4\times 5$ matrix multiplied by a $5\times 5$ matrix from $76$ to $73$, and that from $95$ to $94$ for a $5\times 5$ matrix multiplied by another $5\times 5$ matrix. These results are obtained in characteristic two.
Bilinear complexity, Strassen’s algorithm, Tensor rank
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Computing methodologies Algebraic algorithm
1. Introduction
@@
Efficient matrix multiplication is crucial for accelerating various computational processes, influencing overall computational speed significantly. Therefore, numerous studies have been undertaken to reduce the computational cost from various perspectives. The groundbreaking work of Strassen demonstrated that two $2\times 2$ matrices can be multiplied with only $7$ multiplications (strassen1969gaussian, 9), leading to extensive studies on the complexity of matrix multiplication and the discovery of optimal multiplications for matrices beyond $2\times 2$. For the complexity of matrix multiplication, recent developments continue to aim at finding the upper bound on the matrix multiplication exponent $\omega$. The current best-known bound is $\omega<2.371552$ (williams2023new, 10), improving the previous best bound of $\omega<2.371866$ (duan2022faster, 1). For optimal multiplication methods, notable progress has been made, reducing the multiplication of two $3\times 3$ matrices to $23$ (laderman1976noncommutative, 6), and the multiplication of two $4\times 4$ matrices have been reduced to $49$ using recursively applying Strassen’s algorithm (strassen1969gaussian, 9). Recently, a reinforcement learning approach has further refined the multiplication of two $4\times 4$ matrices to $47$ (Fawzi2022, 2) and improved the best-known result for two $5\times 5$ matrices from $98$ to $96$ multiplications. Subsequently, focusing on the symmetry in matrix multiplications, a topology called flip graph has been introduced (kauers2022flip, 4). By randomly repeating the local transition of vertices over the flip graph, the best-known result for two $5\times 5$ matrices has been improved from $96$ to $95$ multiplications (kauers2022fbhhrbnrssshkalgorithm, 5, 4).
@@
In this study, we first identify two practical issues confronting the flip graph algorithm. The first issue is that the local transition can encounter troublesome “states” which limit further exploration and prevent updating the best number of multiplications. The second issue is that a random search on the flip graph becomes impractical as the matrix size increases within a reasonable computation time. In fact, without setting the results obtained by the reinforcement learning to the initial condition, it would take too long to search from the initial condition and the best number of multiplications would not be reached (kauers2022flip, 4). To address these issues, this study proposes an “adaptive flip graph algorithm”. For the first issue, when encountering troublesome states, the algorithm adaptively transitions vertices in a direction that does not necessarily reduce the number of matrix multiplications, facilitating an escape from the states. We demonstrated that the connectivity of the flip graph is ensured by employing this transition. This proof confirms that all vertices in the flip graph are accessible, guaranteeing comprehensive exploration. For the second issue, instead of conducting a completely random search, the algorithm adaptively constraints the search range, enabling faster exploration. Experiments show that the adaptive flip graph algorithm reduces the number of multiplication for a $4\times 5$ matrix and a $5\times 5$ matrix from $76$ to $73$ and the number of multiplication of two $5\times 5$ matrices from $95$ to $94$, in characteristic 2.
2. Problem Formulation of Matrix Multiplication
@@
Definition 2.1 (Matrix Multiplication Tensor).
@@
(1) $\mathcal{M}_{n,m,p}=\sum_{i,j,k=1}^{n,m,p}a_{i,j}\otimes b_{j,k}\otimes c_{k,i}\in K^{n\times m}\otimes K^{m\times p}\otimes K^{p\times n}.$
@@
Example 2.2 (Standard Classical Matrix Multiplication Algorithm for two $2\times 2$ matrices).
$\mathcal{M}_{2,2,2}$ $=a_{1,1}\otimes b_{1,1}\otimes c_{1,1}+a_{1,2}\otimes b_{2,1}\otimes c_{1,1}$
$+a_{1,1}\otimes b_{1,2}\otimes c_{2,1}+a_{1,2}\otimes b_{2,2}\otimes c_{2,1}$
$+a_{2,1}\otimes b_{1,1}\otimes c_{1,2}+a_{2,2}\otimes b_{2,1}\otimes c_{1,2}$
$+a_{2,1}\otimes b_{1,2}\otimes c_{2,2}+a_{2,2}\otimes b_{2,2}\otimes c_{2,2}.$
@@
Example 2.3 (Strassen’s Algorithm).
@@
In the subsequent discussion, we introduce the following definition for these rank-one tensors.
Definition 2.4 (Matrix Multiplication Scheme).
A finite multiset $S$ of rank-one tensors, whose sum is $\mathcal{M}_{n,m,p}$, is referred to as an $(n,m,p)$-matrix multiplication scheme and $|S|$ is the rank of the scheme.
2.1. Flip Graph Algorithm
A graph, known as the flip graph, has been introduced to represent the transformation relationships between matrix multiplication schemes. A search method on this graph has been proposed to reduce the rank of the scheme (kauers2022flip, 4). This algorithm, which is the basis of the one employed in our study, is reviewed in this subsection.
Definition 2.5 (Flip).
@@ 2.1. Flip Graph Algorithm
Definition 2.6 (Reduction).
For a non-empty set $I\subseteq\{1,2,\ldots,|S|\}$ in a scheme $S$, another scheme $S^{\prime}$ with a rank
$|I|-\text{rank}\left(\sum_{i\in I}\beta^{(r)}\otimes\gamma^{(r)}\right)$ smaller than the rank of S can be constructed using Gaussian elimination if the following conditions hold:
  1. (1)
@@
  2. (2)
    $\text{rank}\left(\sum_{i\in I}\beta^{(r)}\otimes\gamma^{(r)}\right)<|I|$.
@@
Definition 2.7 (Flip Graph).
@@
3:         $S\leftarrow$ Flip of $S$ by Definition 2.5
@@
8:         $S\leftarrow$ Reduction of $S$ by Definition 2.6
@@
3. Limitation and Practical Issues of Flip Graph Algorithm
@@
3.1. Non-Reduction States in Principle
Numerical experiments have revealed a phenomenon where, for the $(4,4,4)$-matrix multiplication scheme, the rank of the scheme does not decrease during the search, even with increasing the number of iterations. This phenomenon is considered to be due to the absence of vertices that can reduce the rank among the vertices accessible by the flip algorithm. Such a state, called a “non-reduction state”, becomes more common as the matrix size increases.
3.2. Practical Non-Reducibility for Large Matrix Multiplication
In random search experiments for the $(5,5,5)$-matrix multiplication scheme, we observed another phenomenon: the rank does not reduce at all from the initial scheme, which is the standard classical matrix multiplication algorithm as shown in Example 2.2. Notably, this phenomenon is not observed for a $4\times 4$ matrix or lower and only appears for order $5$ or larger matrices. This observation implies that as the size of the tensor increases, reaching reducible vertices by the random search becomes more challenging.
The following argument shows that this issue is distinct from the non-reduction states described above. Suppose we take the standard classical matrix multiplication algorithm as in Example 2.2 for the initial scheme. If the matrix is decomposed into submatrices, Strassen’s algorithm as in Example 2.3 can eventually be applied to a $2\times 2$ submatrix, meaning that there is a reducible vertex in the neighborhood of the initial scheme on the flip graph. Thus, this consideration shows that this initial scheme is not in non-reduction states. Even then, it is challenging in a random search to start the rank reduction from the initial scheme.
4. Adaptive Flip Graph Algorithm
@@
4.1. Plus Transition
To address the former issue, we propose the plus transition, which introduces a transition that increases the rank of the scheme. This provides a strategy to escape from the non-reduction state described above, allowing the exploration of schemes that are inaccessible by the flip graph algorithm. The procedure for increasing the rank is defined as a ”Plus” transformation.
Definition 4.1 (Plus).
For a scheme $S$ the $\alpha^{(i)}\neq\alpha^{(j)},\beta^{(i)}\neq\beta^{(j)},\gamma^{(i)}\neq\gamma^{(j)}$ are satisfied, the “Plus” transformation from $S$ to $S^{\prime}$ is defined as follows:
@@ 4.1. Plus Transition
$\rightarrow\alpha^{(i)}\otimes\beta^{(i)}\otimes\gamma^{(i)}+\alpha^{(i)}\otimes\beta^{(j)}\otimes\gamma^{(j)}$
$+(\alpha^{(j)}-\alpha^{(i)})\otimes\beta^{(j)}\otimes\gamma^{(j)}$
$\rightarrow\alpha^{(i)}\otimes(\beta^{(i)}+\beta^{(j)})\otimes\gamma^{(i)}+\alpha^{(i)}\otimes\beta^{(j)}\otimes(\gamma^{(j)}-\gamma^{(i)})$
$+(\alpha^{(j)}-\alpha^{(i)})\otimes\beta^{(j)}\otimes\gamma^{(j)}.$
@@
12:         $S_{t}\leftarrow$ Plus $S_{t}$ by Definition 4.1
@@ 4.2. Edge Constraints
5. Connectivity of flip graphs through plus transitions
In this section, we demonstrate that introducing plus transitions ensures the connectivity of the flip graph for $K=\mathbb{Z}2$, facilitating the exploitation of any matrix multiplication schemes.
Lemma 5.1 ().
In the $(n,m,p)$-matrix multiplication scheme S, it holds that $rank([\alpha^{(1)},\alpha^{(2)},...,\alpha^{(R)}])=n*m$. Similarly, for $\beta$ and $\gamma$, $rank([\beta^{(1)},\beta^{(2)},...,\beta^{(R)}])=m*p$ and $rank([\gamma^{(1)},\gamma^{(2)},...,\gamma^{(R)}])=p*n$.
Proof.
We show that $\text{rank}([\alpha^{(1)},\alpha^{(2)},\ldots,\alpha^{(R)}])=n\times m$. The proof for $\beta$ and $\gamma$ follows a similar argument.
By the definition of matrix multiplication,
(2) $C_{i,j}=\sum_{k=1}^{m}A_{i,k}B_{k,j}$
Considering the definition of the matrix multiplication tensor from Equation 1,
(3) $C_{i,j}$ $=\sum_{r=1}^{R}\gamma_{j,i}^{(r)}\left(\sum_{i_{a},j_{a}=1}^{n,m}\alpha_{i_{a},j_{a}}^{(r)}A_{i_{a},j_{a}}\right)\left(\sum_{i_{b},j_{b}=1}^{m,p}\beta_{i_{b},j_{b}}^{(r)}B_{i_{b},j_{b}}\right)$
From Equation 2, noting that the sum of all terms excluding $B_{k,j}$ is zero, we have
(4) $=\sum_{r=1}^{R}\gamma_{j,i}^{(r)}\left(\sum_{i_{a},j_{a}=1}^{n,m}\alpha_{i_{a},j_{a}}^{(r)}A_{i_{a},j_{a}}\right)\left(\sum_{k=1}^{m}\beta_{k,j}^{(r)}B_{k,j}\right)$
(5) $=\sum_{k=1}^{m}\left(\sum_{r=1}^{R}\beta_{k,j}^{(r)}\gamma_{j,i}^{(r)}\left(\sum_{i_{a},j_{a}=1}^{n,m}\alpha_{i_{a},j_{a}}^{(r)}A_{i_{a},j_{a}}\right)\right)B_{k,j}.$
Comparing this with Equation 2, we find
(6) $A_{i,k}$ $=\sum_{r=1}^{R}\beta_{k,j}^{(r)}\gamma_{j,i}^{(r)}\left(\sum_{i_{a},j_{a}=1}^{n,m}\alpha_{i_{a},j_{a}}^{(r)}A_{i_{a},j_{a}}\right),$
(7) $a_{i,k}$ $=\sum_{r=1}^{R}\beta_{k,j}^{(r)}\gamma_{j,i}^{(r)}\alpha^{(r)}.$
Thus, for any $i,k$, $a_{i,k}$ can be expressed as a linear combination of the vectors $\alpha^{(r)}$. This implies that the matrix constructed by arranging the vectors $\alpha$ for scheme S is of full rank, concluding that $\text{rank}([\alpha^{(1)},\alpha^{(2)},\ldots,\alpha^{(R)}])=n\times m$. ∎
Regarding the scheme transformation by the plus operation as defined in Definition 4.1, the first operation
$\alpha\otimes\beta\otimes\gamma\rightarrow\alpha^{\prime}\otimes\beta\otimes\gamma+(\alpha-\alpha^{\prime})\otimes\beta\otimes\gamma$
is interpreted as splitting the rank-1 tensor $\alpha\otimes\beta\otimes\gamma$ by $\alpha^{\prime}$. In this transition, $\alpha^{\prime}$ is limited to those $\alpha$ existing within the scheme $S$. Under this limitation, we prove that the flip graph is connected on $\mathbb{Z}2$ as follows.
Theorem 5.2 ().
The $(n,m,p)$-flip graph becomes connected over $\mathbb{Z}_{2}$ through the addition of edges by plus transformation.
Proof.
For any two schemes $S$ and $S^{\prime}$, We will show that there is a path from $S$ to $S^{\prime}$. If $|S|<|S^{\prime}|$, $S$ is transformed by applying number of $|S^{\prime}|-|S|$ arbitrary plus transformations. Let $S_{1}$ be the resulting scheme. If $|S|\geq|S^{\prime}|$, let $S_{1}=S$. Next, the elements of $S_{1}$ and $S^{\prime}$ are ordered, and for any given $k$, the $k$-th element of $S_{1}$ is transformed into the $k$-th element of $S^{\prime}$. The elements added by the plus transformation are added after the $k$-th element. According to Lemma 5.1, the $\alpha$ component of the $k$-th element of $S^{\prime}$ can be expressed as a linear combination of $\alpha$ in $S_{1}$. Therefore, by applying the plus transformation repeatedly, the $\alpha$ component of the $k$-th element in $S_{1}$ can be transformed into that of $S^{\prime}$. This process is similarly applied to the $\beta$ and $\gamma$ components, ensuring that the $k$-th element of $S_{1}$ matches that of $S^{\prime}$. Let $S_{2}$ denote this scheme.
For the elements beyond the $|S^{\prime}|$-th position in $S_{2}$, the following transformations are performed. According to Lemma 5.1, for any $i,j$, it is possible to represent $a_{i,j}$ as a linear combination of $\alpha$ in $S_{2}$. Therefore, by repeatedly applying plus transformations, a scheme $S_{3}$ can be constructed such that for any $i,j$, $a_{i,j}\in\{\alpha^{(r)}\}$. Next, using only those $\alpha^{(r)}$ that are in $\{a_{i,j}\}$ for splitting , plus transformations are performed to obtain a scheme $S_{4}$, where $\{\alpha^{(r)}|r>|S^{\prime}|\}\subset\{a_{i,j}\}$. Applying the same operations from $S_{2}$ to $S_{4}$ to $\beta$ and $\gamma$, the scheme $S_{5}$ is obtained.
Finally, in scheme $S_{5}$, for elements beyond the $|S^{\prime}|$-th position, the reductions are repeatedly applied to any pairs of rank-1 tensors. The elements beyond the $|S^{\prime}|$-th position in $S_{5}$ are expressed as $a_{i,j}\otimes b_{k,l}\otimes c_{m,n}$, and their sum is zero. Hence, this operation eliminates all such elements, resulting in $S_{6}=S^{\prime}$. ∎
In this proof, it is proven that the scheme $S$ can be transformed into $S^{\prime}$ through a finite number of transformations. This leads to the following corollary, which appears almost trivial but proves the validity of the basic strategy of the adaptive flip-graph algorithm:
Corollary 5.3 ().
A scheme with the minimal rank can be reached from the initial scheme by a finite number of transitions on the flip graph with edges of plus transformations.
6. Results